An aquarium has a square base with a side lenght of 15 inches. You fill the aquarium with water to a height of 15 inches.
a. Find eh volume of the water in the aquarium.
b. A cubic inch of water weighs approximately 0.036 pound. Find the approximate weight of he wqter in the aquarium.

Answer by Nate(3500)

V = l*w*h
V = 15*15*15 = 3,375 in.^3
Weight = 0.036 * 3,375 = 121.5 pounds